a store mixes brazillian coffee worth $12 per kilogram an venezuelan coffee worth $16 per kilogram the mixture is to sell for $15 per kilogram find out how much of each should be used to make a 216 kilogram mixture.

how many kilogram of the brazillian coffee?
how many kilogram of the venezuelan coffee?

Answer by ewatrrr(24785)


Hi,
Let x represent the kilograms of the brazillian coffee
Then (216kg - x) would represent the kilograms of the venezuelan coffee
question states***
$12/kg*x + $16/kg(216kg-x)= $15/kg*216kg
$1/kg/kg*216kg = $4x
54kg = x, the amount of the brazillian coffee
162kg the amount of the venezuelan coffee (216kg-54kg)
checking our answer
12*54 + 16*162 = 648 + 2592 = 15*216 = 3240